June 15 and 16, 2019

Fishing for the stuck drill

The hook designed as a fishing tool.

Since the drill got stuck at the bottom of the borehole the night before Friday, the drillers have been busy designing and testing a fishing tool with a hook to catch the drill (picture). The tool was welded and assembled Saturday, hard pull tests were made on surface Sunday morning, and the hook descended into the borehole Sunday afternoon. Since then fishing has been ongoing. The hook goes down to touch the top of the drill, and the load of the cable is carefully monitored when pulling up.
If there is no increase in cable load, there has been no luck catching the drill and another fishing attempt is done. At the time of writing (10 PM local), fishing is ongoing.

The water isotope CFA laboratory has had very steady progress over the last weeks. The daily production could almost be predicted already in the morning. Only volcanic ash or very dusty periods have delayed the continuous melting, as they clog the tubing. If tomorrow shows no surprises, the CFA lab will finish the analyses of the samples prepared in 2018. There will thus be a few days break until the science trench starts up next week and new samples will be produced.

Saturday night we all enjoyed a lovely Norwegian-Dutch-Australian-American meal prepared by the MAGPIE team, also known as The Dynamic Four. Later, there was a good party, but due to the drill situation and an early MAGPIE departure Sunday morning, the dome became quiet earlier than usual on Saturdays. Although the weather Saturday evening allowed the ice bar to open, we have had quite some snowfall and snow drift over the weekend. For the skiway, this fresh snow is not ideal considering we are expecting high airport activity the coming week. Hoping for improvement to groom before flights.

What we did Saturday and Sunday:

  1. Designing fishing tool, performing pull tests, and fishing in borehole.
  2. No ice core logging over the weekend.
  3. Station for atmospheric moisture sampling and isotopic measurement is active.
  4. The isotope CFA laboratory bag 3149 to bag 3183 (length 18.70 m).
  5. Sunday MAGPIE picked up 3 stations (SW50km, SW100km, and S50km).
  6. Generator 300h oil check Saturday morning at 7.
  7. First PP measurements carried out in Science trench.
  8. Continuously clearing entrances to trenches due to drift.

EGRIP population is 24.

Weather: Mostly overcast, windy, several cm of fresh snow and drifting snow in camp and on skiway. Temp. -18 °C to -8 °C. Wind: 2 kt to 14 kt turning from SW to N.

FL, Anders Svensson

Lucas is digging his way through the ice sheet in service of the community.

Saturday night a weak pair of sundogs appeared due to light snowfall.